OpenStudy (anonymous):

MEDAL FOR THE BEST ANSWER Of 30 students, 1/3 play sports. Of those who play sports, 2/5 play soccer. How many students play soccer? A. 5 B. 12 C. 10 D. 4

OpenStudy (whpalmer4):

30 students. 1/3 play sports. How many students play sports?

OpenStudy (anonymous):

1/3 of 30 student is 10 ?

OpenStudy (whpalmer4):

Very good. 10 students play sports. 2/5 of them play soccer. How many students play soccer?

OpenStudy (anonymous):

will be 5?

OpenStudy (whpalmer4):

what is \[10*\frac{2}{5}=\]

OpenStudy (whpalmer4):

How many groups of 5 students can you make from 10 students?

OpenStudy (whpalmer4):

2 groups of 5, right? Out of each group of 5 students, 2 of them play soccer. How many students play soccer?

OpenStudy (whpalmer4):

You didn't have any trouble multiplying \[30*\frac{1}{3}\]why is \[10*\frac{2}{5}\]giving you trouble?

OpenStudy (anonymous):

so the answer is 20/5 = 4 ???

OpenStudy (anonymous):

4 is the answer whpalmer4

OpenStudy (whpalmer4):

yes. \[30*\frac{1}{3}*\frac{2}{5} = \frac{30*1*2}{3*5} = \frac{60}{15} = 4\]